FINA Inc.’s assets are $500 million, financed through bank loans, bonds, preferred stocks and common stocks. The amounts are as follows:

Bank loans: $ 100 million borrowed at 9%

Bonds: $180 million, paying 9% coupon with semi-annual payments, and maturity of 5 years. FINA sold its $1,000 par-value bonds for $940 and had to incur $40 flotation cost per bond.

Preferred Stocks: $20 million, paying $15 dividends per share. FINA sold its preferred shares for $210 and had to incur $10/share flotation cost.

Common Stocks: $200 million, the beta of FINA stocks is 1.2, the 90 day Treasury yield is 1%, return on market portfolio is 12 %.

If FINA is subject to 40% tax rate, what is the WACC for FINA?
Answer #1

Answer: Cost of best or Bonds, Kde I (17) + Fus Fts where, 1 - Annual Interest changes t= Tax Rate n = No. of years F = Redeecost of common stocks - Treasury yield + Beta ke ( Return on market poetfolio - Treasu ry yield) = 1 + 1.2 (12-1 = 1 + 13.2 XCost of preferred stacks , kp e D(1 + dt) NS x whose gutest pro D = Preference Dividend dt = Dividend Tax Ns = Net sale proceComputation of Weighted Average cost of capital (Wace Source Amount ($) Weight Cast of Capital overall cost of capiti Bank Lo